### v1

# Noé Arnold, Juan Beccuti, Thunj Chantramonklasri, Matthias Hafner, Nicolas Oderbolz Center for Cryptoeconomics

Abstract
This paper studies how different categories of Ethereum stakers respond to changes in the consensus issuance schedule and how such changes may impact the composition of the staking market. Our findings show that solo stakers are more sensitive to variations in staking rewards compared to ETH holders using centralized exchanges or liquid staking providers. This heightened sensitivity is not explained by the assumptions on cost structure— which would suggest the opposite—but is instead driven by market dynamics. Factors such as superior MEV access and DeFi yields available to other staking methods, as well as the potential presence of inattentive stakers, contribute to this effect. Consequently, a reduction in issuance is likely to crowd out solo stakers while increasing the market share of centralized entities, further concentrating validator power.
I Introduction
As the volume of ETH staked within Ethereum’s Proof-of-Stake (PoS) protocol has grown significantly over the past few years, discussions around the long-term sustainability of the current issuance schedule have become prominent. Central to this debate is the trade-off between the efficient use of validator infrastructure, issuance-driven inflation, and economic security.
A reduction in the consensus issuance schedule has been argued to be important in avoiding an end-game scenario in which most of the circulating supply of ETH is staked. Such an outcome would mean that the supply of staked ETH far exceeds the necessary threshold for economic security, resulting in unnecessary inflationary pressure and inefficient allocation of validator resources. In addition, there are concerns about the potential systemic risks posed by the proliferation of liquid staking tokens and their potential to replace ETH as the de facto currency in the ecosystem. These considerations suggest that it may be prudent to revise the issuance schedule to effectively reduce the incentives for staking.
Conversely, reducing staking rewards may disproportionately impact the profitability of solo stakers, who already represent a small segment of the total validator population. Such a shift could inadvertently accelerate the centralization of Ethereum’s validator ecosystem by driving solo stakers out of the market. In what follows, we aim to test the validity of this argument by examining how different staking methods, and in particular solo stakers, may be affected differently by changes to Ethereum’s consensus issuance policy. Specifically, we focus on answering two key research questions:
  • How do different types of stakers respond to shifts in the issuance schedule?
  • How might these changes impact the decentralization of the Ethereum staking ecosystem?
In the following, we aim to answer these research questions by developing a game-theoretic model of the Ethereum staking market. We run simulations of this model and calibrate it using currently available empirical evidence. Based on the calibrated model, we conduct a comparative analysis of equilibrium outcomes under different consensus issuance schedules.
The empirical literature mostly suggests that solo stakers are more sensitive to changes in consensus issuance yields compared to ETH holders staking through centralized exchanges or liquid staking providers. Consequently, a reduced issuance schedule could be expected to decrease the market share of solo staking while increasing that of centralized exchanges, leading to higher centralization.
Based on our calibrated model, we argue that the cost structures of solo stakers would on their own likely suggest lower supply elasticity with respect to changes in the consensus issuance yield. However, this alone does not fully explain the staking decisions of solo stakers. Instead, their responsiveness is also shaped by market dynamics, where declining issuance intensifies competition for consensus yields. Other staking methods, benefiting from superior MEV access and DeFi yields, are less affected by issuance changes. Strategic solo stakers internalize these competitive pressures, making them more sensitive to shifts in consensus rewards. The results highlights the importance of policy measures like MEV burn in mitigating competitive disadvantages of solo stakers that arise from diminished or highly variable access to MEV revenues.
In addition, we show that reduced issuance disproportionately reduces the profitability of solo staking relative to other staking categories, in part due to the fact that solo staking has fixed costs from the perspective of the ETH holder, while other staking categories do not. While beyond the scope of our model, this reduces the long-term incentives for solo staking and may lead to exits or shifts to solutions such as dSSPs, which can provide stakers with additional DeFi returns and stable MEV revenues.
The structure of this paper is as follows: Section II reviews the existing literature, with a focus on the empirical estimation of the yield elasiticity of staking supply. Section III introduces the game-theoretic model of the Ethereum staking market and outlines its main assumptions and additional model extensions. Section IV reports on the results of the simulation and comparative analysis. Section V discusses the results and Section VI concludes.
II Literature Review
In the following, we focus on research that either estimates the staking behavior of stakers in terms of staking returns or provides empirical evidence on how ETH holders make decisions on which staking method they prefer to use. Several noteworthy contributions have been made to the literature recently.
We turn first to the empirical literature. A recent paper using event studies and empirical elasticity estimation finds that individual stakers in Ethereum exhibit greater sensitivity to relative declines in returns compared to other staking categories (Eloranta and Helminen,, 2025). The authors argue that this suggests that individual solo stakers may be disproportionately affected by changes in network incentives.
Another recent contribution to the literature is based on survey data and semi-structured interviews (Zhu et al.,, 2025). Here, the authors find that solo stakers exhibit heightened sensitivity to changes in the issuance curve relative to other stakeholders, a finding consistent with previous findings. The authors also find some evidence of market segmentation. Large crypto holders tend to view staking rewards as a secondary benefit, with their primary focus being on ETH price appreciation. These individuals consequently tend to prefer liquid staking solutions. On the other hand, retail investors tend to prioritize convenience and brand trust over short-term price fluctuations or minor yield fluctuations. As a result, they often stake through centralized platforms such as Coinbase and are relatively insensitive to fluctuations in staking returns. Finally, institutional stakers tend to emphasize security and regulatory compliance in their staking decisions, with their primary motivation being to protect their ETH holdings from dilution.
Overall, the available empirical literature provides several key insights. First, solo stakers tend to be more sensitive to fluctuations in staking returns than other staking categories. This finding challenges the conventional wisdom that solo stakers are primarily motivated by altruism and a commitment to Ethereum’s decentralization, rather than financial incentives. Furthermore, the literature suggests that the Ethereum staking market is to some extent segmented based on user preferences. In particular, institutions and retail investors tend to focus on the staking services of centralized exchanges such as Coinbase, while DeFi-native stakers tend to prefer liquid staking solutions such as Lido.
In terms of formally describing the Ethereum staking market, some preliminary research has been presented in which the staking allocation decision of ETH holders across different staking methods is modeled as a function of both monetary returns and non-monetary preferences, such as convenience, trust, and commitment to decentralization (Ma,, 2024). The model predicts that changes in the issuance schedule do not affect an individual’s optimal choice of staking method. Contrary to the empirical evidence, the result is that the total market share of a given staking method remains unchanged when the issuance schedule is adjusted. This result depends on the assumption of linear cost functions and, importantly, does not model strategic interactions among stakers competing for issuance returns. In the following, we extend this research by relaxing the assumption of linear cost functions and by modeling strategic interactions between market participants.
III Theoretical Model
This section presents a game-theoretic model incorporating heterogeneous ETH holders and multiple staking alternatives. It outlines key simplifying assumptions and explores potential model extensions. The model framework facilitates analysis of how differing levels of staker sophistication and SSP fee structures influence staking participation and the overall reward structure.
III-A Baseline Model
Building on the empirical evidence reviewed in Section II, the model differentiates ETH holders based on their preferences and level of technical sophistication, which in turn determines the staking options available to them. Specifically, three distinct types of ETH holders are considered:
[…]
Each category consists of a fixed number of players, denoted as $N_{r}$, $N_{t}$, and $N_{ss}$ for Retailers, Techies, and Experts, respectively. For analytical tractability, we assume homogeneity within each group.
In addition, the model includes several staking methods that differ both in the costs they impose on the ETH holder and in the revenue streams they provide. These methods include solo staking, staking through a centralized service provider (cSSP) such as a centralized exchange, and staking through a decentralized service provider (dSSP) that provides additional DeFi revenues through an LST (e.g., Lido).
Consistent with the evidence presented in Section II, we assume a segregated staking market. This means that ETH holders of the Retailer type can either stake with a cSSP or opt out of staking altogether, arguably reflecting their preference for the convenience of centralized exchanges or their limited technical expertise that prevents them from engaging in solo staking or using a dSSP. Techies choose to stake through a dSSP or not to stake at all. This is consistent with Zhu et al., (2025), as they are assumed to value liquidity and prefer decentralized intermediaries that provide additional returns through liquid staking tokens (LSTs). Finally, Experts engage exclusively in solo staking or abstain from staking, reflecting their strong preference for decentralization, which precludes their reliance on any intermediaries. In the following, we will use the terms solo staker and expert interchangeably.
Depending on the staking option they choose, ETH holders may receive revenues from three different sources: 1) consensus layer revenue, 2) MEV revenue, and 3) additional DeFi yield.
Consensus layer revenues are determined by the protocol’s issuance schedule. The annual consensus yield provided to an ETH holder when staking is denoted as $y_{i}(D)$ and under the current issuance schedule is given by
$y_{i}(D)=\frac{cF}{\sqrt{D}}=\frac{2.6\cdot 64}{\sqrt{D}},$
where $D$ represents the total amount of staked ETH staked. In the following, we will essentially compare model results under the current issuance schedule $y_{i}(D)$ with alternative specifications $y_{i}^{\prime}(D)$.
The second revenue stream is MEV revenue. Validators selected as proposers in block production can generate additional revenue from structuring transactions. Let $N=D/32$ denote the total number of validators, the probability that an ETH holder controlling $d_{i}/32$ validators will be selected as a proposer is $d_{i}/D$. Conversely, if a staker joins a staking pool (i.e., stakes with an intermediary), proposers can share MEV revenue with validators. The probability of any validator in the pool being selected as a proposer is $d_{pool}/D$, and a validator with stake $d_{i}$ receives a share of $d_{i}/d_{pool}$. Thus, the expected MEV revenue for a stakeholder in a pool remains the same as for a solo staker. In both cases, the expected annual MEV revenue $y_{v}$ for an ETH holder $i$ is theoretically given by
$y_{v}\cdot\frac{d_{i}}{D}.$
The third source of income is additional DeFi income. ETH holders using a dSSP can earn additional income by reinvesting the LST they receive. The annual return from this income source is assumed to be exogenous and is simply defined as $y_{d}$.
Staking entails costs for ETH holders, which vary depending on the staking solution. In this model, each staking method is associated with a distinct cost function that may include fixed and nonlinear variable costs
$C(d)=C+c\,d^{\alpha}.$
In addition, ETH holders who stake through an intermediary incur fees, which are modeled in the following as a tax on staking revenues and as such are treated separately from the cost functions. The fee rate for the cSSP is denoted as $f_{c}$ and the fee rate for the dSSP is denoted as $f_{d}$, both of which are assumed to be exogenous in the baseline model.
Under the assumption of a constant ETH price, thus isolating staking decisions from price volatility and inflation effects, each staker maximizes expected annual profit. Experts choose a deposit $\hat{d}_{ss}\geq 0$ to maximize
$\max_{\hat{d}_{ss}}\,\Big{\{}y_{i}(D)\cdot\hat{d}_{ss}+y_{v}\cdot P_{ss}(D)-\Big{(}C_{ss}+c_{ss}\cdot\hat{d}_{ss}^{\alpha_{ss}}\Big{)}\Big{\}},$ (1)
where
$D=\hat{d}_{ss}+(N_{ss}-1)\cdot d_{ss}+N_{r}\cdot d_{r}+N_{t}\cdot d_{t}\in[0,120M]$
and $P_{ss}(D)$ is the probability that a Solo Staker choosing $\hat{d}_{ss}$ becomes a proposer
$P_{ss}(D)=\frac{\hat{d}_{ss}}{\hat{d}_{ss}+(N_{ss}-1)\cdot d_{ss}+N_{r}\cdot d_{r}+N_{t}\cdot d_{t}}$
Assuming homogeneity, all solo stakers choose the same deposit in equilibrium, i.e., $\hat{d}_{ss}^{*}=d_{ss}$.
Techies choose a deposit $\hat{d}_{t}\geq 0$ to maximize
$\max_{\hat{d}_{t}}\,\Big{\{}(1-f_{d})\cdot\Big{[}y_{i}(D)\cdot\hat{d}_{t}+y_{v}\cdot P_{t}(D)\Big{]}+y_{d}\cdot\hat{d}_{t}$
$-\Big{(}C_{t}+c_{t}\cdot\hat{d}_{t}^{\alpha_{t}}\Big{)}\Big{\}},$ (2)
with
$P_{t}(D)=\frac{\hat{d}_{t}}{N_{ss}\cdot d_{ss}+\hat{d}_{t}+(N_{t}-1)\cdot d_{t}+N_{r}\cdot d_{r}}.$
Homogeneity again implies $\hat{d}_{t}^{*}=d_{t}$ in equilibrium.
Finally, Retailers choose a deposit $\hat{d}_{r}\geq 0$ to maximize
$\max_{\hat{d}_{r}}\,\Big{\{}(1-f_{c})\cdot\Big{[}y_{i}(D)\cdot\hat{d}_{r}+y_{v}\cdot P_{r}(D)\Big{]}$
$-\Big{(}C_{r}+c_{r}\cdot\hat{d}_{r}^{\alpha_{r}}\Big{)}\Big{\}},$ (3)
with
$P_{r}(D)=\frac{\hat{d}_{r}}{N_{ss}\cdot d_{ss}+N_{t}\cdot d_{t}+\hat{d}_{r}+(N_{r}-1)\cdot d_{r}}.$
ETH holders act strategically, anticipating the staking decisions of other market participants when determining their own profit-maximizing level of stake. The resulting model solution thus represents a Nash-equilibrium. We refer to Section -A of the Appendix for a detailed derivation.
III-B Model Extensions
We extend the baseline game theoretic model by sequentially introducing important refinements to the characteristics of stakers. First, we introduce solo stakers who face a larger variance in MEV payoffs than other stakers. Next, we relax the assumption that all stakers are fully strategic by introducing inattentive agents who do not respond to changes in the issuance curve. Finally, we introduce an intermediary dSSP with market power.
III-B1 Variability in MEV rewards for solo stakers
In the baseline model, expected MEV revenues are assumed to be identical whether an ETH holder chooses solo staking or joins a pool through an SSP. However, since MEV revenues can vary substantially over time, this formulation does not account for the fact that solo stakers may take several years to propose a block with high MEV, unlike stakers in a pool who can benefit from smoothing MEV revenues over time. To account for such a difference, we introduce the variance over MEV revenues for a (moderately risk-averse) solo staker, while assuming that this variance can be neglected for the other types, i.e., that pools are large enough. Thus, the maximization problem of a solo staker (assumed to be of type Expert) becomes [^1] See Section -B of the Appendix for the derivation.
$\max_{\hat{d}_{e}}\quad y_{i}(D)\cdot\hat{d}_{e}+[y_{v}\cdot P_{ss}(D)-y_{v}^{2}\cdot P_{ss}(D)^{2}]$
$-(C_{ss}+c_{ss}\cdot\hat{d}_{e}^{\alpha_{e}}),$ (4)
when we use
$E[Y_{MEV}]-Var[Y_{MEV}]$ $=y_{v}\cdot P_{ss}(D)-y_{v}^{2}\cdot P_{ss}(D)^{2}.$
III-B2 Inattentive Agents
Further, we divide the Retailer category into two subgroups: (1) institutions and (2) inattentive retail investors. The latter group is assumed to be unresponsive to changes in the issuance yield. This behavior could be interpreted as the agents being inattentive to changes in issuance or that their profit-maximizing decisions result in a corner solution, as is the case in the theoretical results obtained by Ma, (2024). The primary purpose of incorporating this type of agent into the model is to assess how their presence influences the staking behavior of other types.
III-B3 Intermediary dSSP with Market Power
So far, we have simplified the role of decentralized staking service providers (dSSPs) in the market. In particular, the above model specifications implicitly assume that dSSPs operate under perfect competition and that they provide the staking service themselves.
A more realistic approach would be to model dSSPs as intermediaries or middleware between stakers and centralized staking service providers (cSSPs). For example, if an ETH holder decides to stake with Lido, the stake is managed by a cSSP. Lido itself does not directly provide the staking service, but rather facilitates the connection between the ETH holder and the cSSP. In this arrangement, the ETH holder benefits from Lido by receiving a derivative representing its stake. Figure 1 illustrates this setup in a simplified way.
Figure 1: dSSP as Intermediaries between ETH holders and cSSPs
We extend the model to capture such interactions between an intermediary, stakers, and multiple centralized staking service providers (cSSPs). In doing so, our framework specifically captures Lido, the dominant player in this intermediation space. Instead of assuming perfect competition, we model the decentralized staking service provider (dSSP) as a player with market power. We incorporate this extension in addition to the previously described extensions that account for inattentive retail stakers and the higher variance in execution rewards faced by solo stakers.
The goal is to examine how changes in the issuance curve affect the dSSP’s fee-setting decisions and, consequently, overall staking levels. We maintain the timing structure from previous model iterations: the dSSP sets its fees first, followed by the stakers setting their deposit levels.
In the following, we describe the updated profit maximizing behaviors of the players shown in Figure 1: the intermediary dSSP (Lido), its potential users (ETH holders of type Techie), and the cSSPs. The strategic behavior of the other actors, in particular the ETH holder types Expert and Retailer, remains unchanged.
We maintain the assumption of a segregated staking market and assume that the ETH holder type Techie has access to choose the dSSP as an intermediary. In doing so, they are indifferent to the specifics of the dSSP-cSSP relationship; their decision is based solely on Lido’s fees ($f_{d}$), their expected rewards, and associated costs. Essentially, they solve the same optimization problem as in equation (2): each Techie selects $\hat{d}_{t}\geq 0$ to maximize their profit
$\max_{\hat{d}_{t}}\Big{\{}(1-f_{d})\cdot\left[y_{i}(D)\cdot\hat{d}_{t}+y_{v}\cdot P_{t}(D)\right]$
$+y_{d}\cdot\hat{d}_{t}-(C_{t}+c_{t}\cdot\hat{d}_{t}^{\alpha_{t}})\Big{\}},$ (5)
To maximize its profits, the dSSP (Lido) sets both the fee $f_{d}$ charged to Techies and the fee $\hat{f}_{d}$ paid to cSSPs. The dSSP holds the bargaining power to determine these fees, as there are many cSSPs competing to be integrated into the platform. We assume that the dSSP charges a uniform fee to each cSSP and distributes stakes evenly among them. The profit maximization problem of the dSSP is therefore:
$\max_{f_{d},\hat{f}_{d}}$$\Big{\{}(f_{d}-\hat{f}_{d})\cdot$
$\left[y(D_{-t},D_{t}(f_{d}))\cdot D_{t}(f_{d})+y_{v}(D_{-t},D_{t}(f_{d}))\right]$
$-C_{i}(D_{t}(f_{d}))\Big{\}},$ (6)
[…]
where $D_{t}=\sum\hat{d}_{t}$, and $D_{-t}$ denotes the stakes from non-Techies. We assume that Lido enjoys economies of scale, i.e., $C_{i}(D_{t})$ is a concave function. In setting $\hat{f}_{d}$, Lido ensures that cSSP profits, $\pi_{cSSP}$, remain at or above zero. In other words, Lido considers the cSSPs’ participation constraints (PC) to secure sufficient cSSP participation for allocating Techies’ stakes.
Consider any cSSP $m=1...M$. This cSSP sets the fee $f_{c}^{m}$ charged to stakers who bypass the dSSP (Lido) and stake directly with it (see Staker 1 in Figure 1).
$\max_{f_{c}^{m}}\quad f_{c}^{m}\cdot\left[y(D)\cdot D_{i}^{m}+y_{v}(D_{i}^{m},D)\right]$
$+\hat{f}_{d}\cdot\left[y(D)\cdot D_{t}^{m}+y_{v}(D_{t}^{m},D)\right]$
$-C_{c}(D_{t}^{m}+D_{i}^{m}),$ (7)
where $D_{i}^{m}$ represents the total stakes from institutions that choose to stake with $m$. As previously mentioned, we assume that Lido distributes $D_{t}$ uniformly across all the cSSPs:
$D_{t}^{m}=\frac{D_{t}}{M}.$
Again, we assume that cSSPs benefit from economies of scale, i.e., $C_{c}(D_{t})$ is a concave function.
IV Results
In the following, we compare the equilibrium outcomes of the game-theoretic model under two different issuance schedules. The first issuance schedule corresponds to the one currently implemented by the Ethereum protocol:
$y_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}},$
where $D$ represents the staking deposit level. The second issuance schedule follows one of the alternative proposals presented by Elowsson, (2024):
$y^{\prime}_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}\cdot(1+k\cdot D)},$
where $k=2^{-25}$ and $D$ again represents the total amount of ETH staked.
IV-A Model Simulation
As a first step, the general relationships between the model parameters and the resulting equilibrium outcomes are explored. The goal is to develop an intuition for the relevant factors driving equilibrium staking decisions and how they might drive differences in the outcomes under the two consensus issuance schedules. This is done by running simulations of the model, randomly varying the parameter values in each simulation run. We conduct 5 million simulation runs for each of the two issuance schedules.
The simulations yield several key observations that offer insight into how different model parameters influence changes in equilibrium outcomes when transitioning from $y_{i}(D)$ to $y^{\prime}_{i}(D)$:
  • Observation 1: Higher marginal costs lead to smaller adjustments in the equilibrium staking supply when transitioning from $y_{i}(D)$ to $y^{\prime}_{i}(D)$. This effect is illustrated in Figures 3(a) - 3(c), using solo staking as an example.
  • Observation 2: When stakers have access to additional MEV revenues or DeFi yields, they exhibit a weaker response in their equilibrium staking supply when transitioning from $y_{i}(D)$ to $y^{\prime}_{i}(D)$. Figures 3(d) and 3(e) illustrate this effect using staking via dSSP as an example.
  • Observation 3: The sensitivity of one type of agent to changes in the issuance schedule affects the sensitivity of other agent types to changes in the issuance schedule. For instance, when Techies become less sensitive to issuance schedule changes—such as when additional DeFi yields are high—the sensitivity of Experts to a change in the issuance schedule increases. Figure 3(f) illustrates this effect.
Figure 2: Relationship between cost parameters and change in equilibrium staking decision
(a) $D_{ss}$ vs. $C_{ss}$
(b) $D_{ss}$ vs. $c_{ss}$
(c) $D_{ss}$ vs. $\alpha_{ss}$
(d) $D_{dSSP}$ vs. $y_{v}$
(e) $D_{dSSP}$ vs. $y_{d}$
(f) $D_{ss}$ vs. $y_{d}$
Note: The plots display the binned average percentage change in the total staking supply of different staking methods when transitioning from $y_{i}(D)$ to $y^{\prime}_{i}(D)$, categorized by different levels of model parameters. The number of bins is set to 20.
Recall that the empirical evidence discussed in Section II suggests that solo stakers tend to be more sensitive to changes in staking yield than other staking categories. Within the framework of the model, an increased sensitivity of solo stakers could arise from several factors. One possible explanation is that the marginal cost of solo staking is lower relative to other staking methods, making solo stakers more responsive to fluctuations in rewards. Alternatively, solo stakers may have more limited access to external revenue opportunities or may discount expected MEV revenues more than other market participants due to their variability. Finally, since ETH holders compete for consensus issuance yields and we assume strategic solo stakers, either of these dynamics could be exacerbated by other types of ETH holders being relatively increasingly less elastic to changes in consensus issue yields.
IV-B Comparative Analysis
Below we analyze the equilibrium results of the game theoretic model based on specific parameter calibrations. The calibration of the model parameters, in particular the underlying assumptions about cost structures and revenue streams, is described in detail in Section -C of the Appendix. In essence, the calibration approach relies on setting cost parameters such that, under the current issuance schedule, the model produces today’s distribution of stake offerings. However, the calibrated model may not reflect the true population parameters, especially if the model does not capture all of the factors that drive agents’ decisions in the staking market.
IV-B1 Baseline Model
We first compare the equilibrium staking outcomes of the calibrated baseline model with the equilibrium outcomes of the calibrated model under the alternative issuance schedule $y^{\prime}_{i}(D)$. Table I presents the total amount of ETH staked and the corresponding profits for each type of ETH holder under various issuance schedules, as well as the relative changes in these values resulting from the change in the issuance schedule.
TABLE I: Nash Equilibria of Baseline Model
Type Num $y_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}}$ $y^{\prime}_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}\cdot(1+k\cdot D)}$ $\Delta$
Deps Ratio Profita Deps Ratio Profita Deps Profit
e 25K 0.9M 2.7% 0.8% 0.74M 2.97% 0.2% -17.8% -75%
t 200K 18M 54.1% 1.8% 14.3M 57.5% 1.6% -20.6% -11.1%
r 925K 14.4M 43.2% 0.95% 9.8M 39.53% 0.78% -31.5% -17.9%
Total - 33.3M - - 24.84M - - -25.4% -
a Profit per ETH staked.
The first observation to make is that a change in the issuance schedule results in an approximate 25% reduction in the total amount of staked ETH. The relative decrease is most pronounced for the ETH holder type Retailer, followed by Techie and Expert. Consequently, although the overall staking supply declines across all types, the market share of solo staking and staking via dSSPs increases. This result may be explained by Experts facing high marginal costs when solo staking, making them less sensitive to changes in the issuance schedule, as well as Techies having access to additional external DeFi yields, again making them less sensitive to changes in the issuance schedule. Despite a substantial reduction in their equilibrium staking supply, all ETH holder types experience lower staking profits per ETH staked under the adjusted issuance schedule. This decline is particularly pronounced for the Expert type, who uniquely must cover their fixed costs and higher average costs with a reduced revenue base.
IV-B2 Extended Model - Expected MEV Revenues
Next, we repeat the above analysis for the extended version of the model, where the type Expert is assumed to be risk-averse and as such discounts future MEV revenues. Table II again reports the resulting equilibrium outcomes under the two issuance schedules.
TABLE II: Nash Equilibria of Extended Model with Discounted MEV Revenues
Type Num $y_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}}$ $y^{\prime}_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}\cdot(1+k\cdot D)}$ $\Delta$
Deps Ratio Profita Deps Ratio Profita Deps Profit
e 25K 0.77M 2.3% 0.6% 0.58M 2.3% -0.01% -24.6% -100%
t 200K 18M 54.3% 1.8% 14.3M 57.8% 1.6% -20.6% -11.1%
r 925K 14.4M 43.4% 0.95% 9.9M 39.85% 0.78% -31.25% -17.9%
Total - 33.2M - - 24.76M - - -25.4% -
a Profit per ETH staked.
When incorporating the variance in MEV rewards for a risk-averse and small solo staker of type Expert, their relative reduction in staking supply becomes more pronounced compared to the baseline model. As outlined in Section IV-A, this outcome is driven by the following dynamic: if agents expect to have access to lower expected MEV revenues, they become more sensitive to changes in the issuance schedule. As a result, a change in the issuance schedule no longer leads to an increase in the market share of solo stakers. At the same time the staking outcomes for the remain staking solutions and ETH holder types remain virtually unchanged relative to the baseline model. Since Expert solo stakers represent only a small fraction of the overall population, their heightened sensitivity to issuance changes has little impact on Techies and Retailers. Additionally, the negative effect of issuance reductions on the expected profits per staked ETH for solo stakers is further exacerbated, with expected profits even turning negative.
IV-B3 Extended Model - Inattentive Agents
Now, we incorporate to the model with the higher variability in MEV rewards for solo stakers, the differentiation of the ETH holder type Retailer into (1) institutions and (2) an inattentive retail investor, as described in Section III-B2. Table III reports the corresponding results.
TABLE III: Nash Equilibria of Extended Model with Inattentive Agents
Type Num $y_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}}$ $y^{\prime}_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}\cdot(1+k\cdot D)}$ $\Delta$
Deps Ratio Profita Deps Ratio Profita Deps Profit
e 25K 0.78M 2.4% 0.6% 0.57M 2.2% -0.0% -26.9% -100%
t 200K 18.1M 54.7% 1.8% 13.7M 53.3% 1.6% -24.3% -11.1%
i 300K 7.7M 23.3% 0.95% 4.9M 19.1% 0.76% -36.4% -20.0%
r NA 6.5M 19.6% NA 6.5M 25.3% NA 0% NA
Total - 33.1M - - 25.7M - - -22.4% -
a Profit per ETH staked.
In the presence of the inattentive ETH holder type—whose staking supply remains constant across both issuance schedules—the relative reduction in the overall staking supply is somewhat lower than in the baseline model. Meanwhile, the ETH holder types that do adjust their staking supply in response to the issuance schedule experience a decline in their market shares under the adjusted schedule. Again this is consistent with the intuitions developed in Section IV-A. When one ETH holder type does not modify its staking supply in response to an issuance reduction, the adjustments made by the other types tend to be more pronounced. The solo staking type Expert still experiences a larger reduction in their staking supply than the Techie type, who continues to benefit from external DeFi yields, making their equilibrium staking supply less sensitive to changes in protocol issuance. Consequently, the decline in staking profits per ETH staked is even greater for the Expert type.
IV-B4 Extended Model - Intermediary dSSP with Market Power
Finally, we conduct a comparative analysis of equilibrium outcomes under the assumption of an intermediary dSSP, similar to Lido, that acts as a platform between ETH holders and SSPs. Importantly, the fee set by the dSSP is endogenous to the model and therefore subject to the strategic response of the dSSP to a change in the issuance schedule. In solving the model, we make some additional assumptions and simplifications relative to the model presented in Section III-B3:
  • We assume that, as a result of change in the consensus issuance schedule, the dSSP (Lido) may change $f_{d}$ but not $\hat{f}_{d}$.
  • To simplify our analysis, we do not explicitly solve the dSSP’s full profit maximization problem as outlined above, since our primary focus is on determining the direction in which the fee should move. To solve exactly for the optimal fee, we would first need to calibrate Lido’s cost function. Instead, given a cost structure with economies of scale, we calculate Lido’s profits under the current issuance schedule and the fee assumed in the baseline model calibration. Then, under the updated issuance yield curve, we simply identify the new fee as the fee that would yield a similar profit, and then use that fee to derive the new staking equilibrium.
  • Similarly, we do not solve for the optimal choice of $f_{c}^{m}$ for the cSSPs, and when comparing equilibrium outcomes under different issuance schedules, we simply consider the variable to be fixed to the same calibration as in the baseline model. This may cause $f_{c}^{m}$ to be set too high, since in the baseline version of the model, cSSPs are assumed to be more like a CeX, and in the current extension, they are more akin to SSPs like P2P and Figment. However, this assumption ensures a better comparability with the previous results and we are therefore better able to isolate the effect of the market power of the dSSPs on the equilibrium results.
Table IV reports the results of the comparative analysis under these assumptions, in particular the fees set by the dSSP and the distribution of stakes across different ETH holder types and staking methods.
TABLE IV: Nash Equilibria of Extended Model with Intermediary dSSP
Type Num $y_{i}(D);f_{d}=10\%$ $y^{\prime}_{i}(D);f_{d}=13\%$ $\Delta$
Deps Ratio Deps Ratio Deps
ss 25K 0.78M 2.4% 0.57M 2.2% -26.9%
t 200K 18M 54.5% 13.4M 52.5% -25.5%
i 300K 7.7M 23.3% 5M 19.6% -35%
r NA 6.5M 19.7% 6.5M 25.5% 0%
Total 33M 25.5M -22.7%
a Dep = Deposits, Ratio = Staking Ratio, $\Delta$ = Change in Deposits
The adjustment of the issuance schedule is associated with an increase of $30\%$ in $f_{d}$. As a result, the decrease in the total amount staked through the dSSP is greater and the staking distribution becomes more concentrated among the cSSPs compared to the previous model iterations where the dSSP fees did not adjust in response to a change in the issuance schedule. Put differently, the increase in $f_{d}$ primarily impacts stakes from Techies, who reduce their participation. Institutions, in turn, benefit from this reduced Techie participation, increasing their staking level by $2\%$—approximately equivalent to the decrease in Techie participation. Since cSSPs are the providers for institutional stakes, this shift could be interpreted as an increase in concentration within the staking market.
V Discussion
It is important to note that the modeling approach above does not consider entry or exit in the sense that we assume a fixed number of agents per ETH holder type. Further, the model does not allow all agents to switch between different staking methods. Based on these assumptions, the model framework may be able to model the short-run dynamics of the Ethereum staking market, but may fall short when considering the long-run perspective. Consequently, the results presented here reflect a short-term perspective on how the staking market might respond to changes in the issuance schedule. In the long run, the relative profitability of different staking solutions is likely to influence market shares, as new entrants or existing market participants opt for profitable staking methods and existing stakers exit unprofitable ones. In the following, we therefore frame the discussion of the above results within a short-term perspective.
V-A Short-term Perspective
The results of the calibrated game theoretic model, where solo stakers are modeled as experiencing disutility from the variability of MEV revenues and where a fraction of individuals staking through centralized providers are assumed to be inattentive, are broadly consistent with the empirical evidence suggesting that solo stakers have a higher yield elasticity of staking supply. Under these assumptions, the staking supply of solo stakers is more responsive to changes in the issuance schedule than that of other types of stakers.
This result underscores the fact that examining cost structures alone is not enough to explain the behavior of staking market participants. Even with cost structures that would otherwise make solo stakers inelastic to changes in consensus yields, the competitive environment in the broader staking market can still make solo stakers vulnerable to being crowded out of the staking market when consensus yields fall. This is because even in a segregated staking market where existing solo stakers do not switch to other staking solutions at all, other market participants may be relatively insensitive to changes in consensus issuance and thus put pressure on solo stakers to reduce their staking supply when consensus yields fall. This result also has important implications for the broader economic policy mix of the Ethereum protocol. It suggests that policies that help improve the competitive standing of solo stakers relative to other staking methods, for example the implementation of MEV burn, could help mitigate the otherwise negative effects of an issuance reduction on the market share of solo stakers.
Assuming that tee model specification with all extensions most closely reflects empirical reality, the market share of solo stakers is expected to decrease under the proposed issuance schedule, while the market share of centralized exchanges would increase. This suggests that a reduction in issuance could help limit the market share of liquid staking providers and the resulting dominance of LSTs within the Ethereum ecosystem, at least in the short run. At the same time, a reduction in issuance may come at the cost of increased market share for centralized providers. Combined with a declining market share of solo stakers, this could contribute to an overall higher degree of centralization in the Ethereum validator set.
V-B Long-term Perspective
Across all model specifications evaluated above, a change in the issuance schedule is associated with a disproportionately negative impact on the expected profitability of solo staking relative to other staking methods. Over time, this short-run decline in profitability may increase incentives for solo stakers to either exit the market or transition to alternative staking methods. On the other hand, the profitability of Techies staking via dSSPs that offer additional DeFi yields is least affected by the decline in issuance, particularly also in comparison to staking via centralized exchanges. In the long run, individual LSTs (or LRTs) may therefore continue to gain market share. The long-term effects on the centralization of the Ethereum validator set are thus ambiguous. Based on the results above, both the most decentralized staking option (solo staking) and the most centralized option (staking via CEXs) become less attractive in the long run.
However, exits and market participants switching between staking methods as a result of diminished profitability are not the only factors that could lead to diminished solo staking and an increasing dominance of liquid staking solutions in the market. This is because, over time, the divergence in profits between the different staking methods leads to a redistributive effect between the different staking methods. If stakers are able to reinvest their profits from staking, and staking profits are comparatively low for solo staking, their market share will continue to diminish. To illustrate this effect, we use the Nash equilibria of the baseline model outlined in Table I as a starting point and chart the cumulative changes in staking levels and staking ratios over time, assuming stakers reinvest their profits (revenues minus costs) and do not reevaluate their optimal level of staking supply over time. The results are shown in Figure 3. Figure 3 depicts the cumulative percentage change in the total staking level relative to the starting point. Similarly, Figure 4 depicts the cumulative change in the share of total ETH supply being staked. As such, Figure 4 also takes into account the effects of dilution.
This thought experiment illustrates the potential redistributive effects arising from differences in profitability, ultimately leading to the gradual displacement of solo stakers from the market over the long term. Moreover, the calculations show that a reduction in the consensus issuance schedule noticeably accelerates this process, as the profitability of solo stakers is disproportionally affected. How strong this effect is in reality depends to a large degree on the extent to which stakers actually reinvest their staking profits. Anecdotal evidence suggests that ETH holders who stake through intermediaries are significantly better positioned to reinvest profits compared to small solo stakers. Given that the protocol sets a maximum effective balance for validators, smaller solo stakers encounter practical difficulties in expanding their revenue base through the reinvestment of staking rewards. In contrast, ETH holders utilizing intermediaries can more effectively compound their returns due to the ability to pool staking profits collectively.
Figure 3: Cumul. long-run effects of profit reinvestment (staking level)
Figure 4: Cumulative long-run effects of profit reinvestment (staking ratio)
Note: Full line represents current issuance yield; dashed line represents adjusted issuance yield; Blue line represents Techies; Red represents Retailers; Yellow represents solo stakers.
VI Conclusion
We show that solo stakers are more responsive to changes in consensus issuance yields than ETH holders staking through centralized exchanges or liquid staking providers. As issuance declines, solo staking is expected to lose market share, while centralized exchanges gain dominance.
This heightened sensitivity of solo stakers is not solely explained by their cost structures but is instead driven by the competitive dynamics of the staking market. Other staking methods benefit from superior MEV access and DeFi yields, making them less responsive to issuance changes and creating a crowding-out effect on solo stakers. These findings underscore the importance of mechanisms such as MEV burn in mitigating the negative impact of issuance reductions on solo stakers.
Additionally, reduced issuance disproportionately lowers solo staking profitability and hence the incentives for solo staking which may drive exits or shifts toward alternative staking solutions like dSSPs with DeFi yields over time.
The present analysis has certain limitations, like the assumptions used in calibrating the model. While these assumptions align with existing literature, they remain open to debate. Future research should prioritize refining the estimation of cost structures across different staking methods, which could enhance the accuracy and robustness of the model framework.
Still, we expect that the model developed here serves as a foundation for further research and extensions, providing a flexible framework for analyzing staking dynamics under various protocol adjustments. For tractability and compatibility with the broader model, we introduced several simplifying assumptions in the model extension with endogenous dSSP fees. However, we believe that our model could be expanded to facilitate a deeper exploration of the competitive dynamics between staking intermediaries.